how would you do this if the exponent is a negative such as something that looks like this 64^-2/3?
Answer by lwsshak3(11628)

how would you do this if the exponent is a negative such as something that looks like this 64^-2/3?
**
There is a rule in algebra which says if you move a term with an exponent from numerator to denominator or vise-versa, the sign of the exponent reverses.
64^-2/3=1/64^2/3=1/16